OpenStudy (31356):

The velocity of a baseball after it is dropped from the top of a building. Is this linear or non linear? Medal for best answer!

OpenStudy (31356):

@Abdulhameed

OpenStudy (31356):

The slope of the line y= 2x-1 is multipled by 2 and the y intercept is shifted down four units. What is the equation of the new line?

OpenStudy (anonymous):

linear

OpenStudy (anonymous):

Y=mx+b So for the first line y=2x-1 M=2 b=-1

OpenStudy (31356):

Okay thanks! :)(:

OpenStudy (anonymous):

M is the slope and b is the y intercept

OpenStudy (31356):

Okay...

OpenStudy (31356):

Can you help me with another question?

OpenStudy (anonymous):

so the slope of the new line =2M=..? and y intercept of the new line =b-4=..?

OpenStudy (31356):

Dylan rented a car from New Car Rental for 7 days and drove it for 387 miles. the rental company charges $24.95 a day plus 8 cents a mile. how much did Dylan have to pay to rent the car?

OpenStudy (anonymous):

the money he should pay = 7(24.95)+0.08(387)=...?

OpenStudy (31356):

I found out the answer :)(:

OpenStudy (31356):

Let me post another question ;)

OpenStudy (anonymous):

okay

OpenStudy (31356):

My question doesn't post for some reason ;(

OpenStudy (31356):

Nathan split 32 ounces of hamburger meat into patties that each weighed 1/4 pound. How many hamburger patties did he have? (1 pound =16 ounces)

OpenStudy (31356):

@Abdulhameed

OpenStudy (anonymous):

An ounce =0.0625 pound right?

OpenStudy (31356):

Trying to remember....

OpenStudy (anonymous):

16 ounces =1 pound 32 ounces = 2 pounds

OpenStudy (anonymous):

each patty is 1/4 pound number of patties =2 pounds /1/4= 2 *4=8 patties